# suppSupplementary Material References

### v1

@@ Abstract
Traditional network analysis focuses on binary edges, while real-world relationships are more nuanced, encompassing cooperation, neutrality, and conflict. The rise of negative edges in social media discussions spurred interest in analyzing signed interactions, especially in polarized debates. However, the vast data generated by digital networks presents challenges for traditional methods like Stochastic Block Models (SBM) and Exponential Family Random Graph Models (ERGM), particularly due to the homogeneity assumption and global dependence, which become increasingly unrealistic as network size grows. To address this, we propose a novel method that combines the strengths of SBM and ERGM while mitigating their weaknesses by incorporating local dependence based on non-overlapping blocks. Our approach involves a two-step process: first, decomposing the network into sub-networks using SBM approximation, and then estimating parameters using ERGM methods. We validate our method on large synthetic networks and apply it to a signed Wikipedia network of thousands of editors. Through the use of local dependence, we find patterns consistent with structural balance theory.
@@ 1 Introduction
Nevertheless, the modeling of signed networks remains challenging. On the one hand, the class of models specifically designed for signed networks is relatively limited, since most models were developed for unsigned networks. On the other hand, the massive size of modern networks poses separate challenges, as even for unsigned networks many models struggle to scale computationally. At the intersection of these two issues, the need to model both signed relations and large networks, there are little to no methods available.
Two of the most prominent models for network analysis are Stochastic Block Models (SBM, Nowicki and and Snijders, 2001) and Exponential Family Random Graph Models (ERGM, Lusher et al., 2013). SBMs are a latent variable model in which all dependence between edges is accounted for by latent group membership. ERGMs, on the other hand, allow for more complex local dependencies through so-called sufficient statistics (De Nicola et al., 2023). Signed network extensions exist for both SBM and ERGM (Jiang, 2015; Fritz et al., 2025) with inherent shortcomings: While the SBM can represent cohesive subgroups among the nodes, it implies conditional independence for forming edges between nodes. This assumption is unrealistic for networks exhibiting more dependence among edges. ERGMs, on the other hand, offer a flexible framework to analyze general network properties. For a parsimonious model, these properties are commonly assumed to be homogeneous across the entire network. The induced global dependence structure, where each edge may technically depend on every other edge in the same way, gives rise to undesirable behavior specifically for large networks (Schweinberger, 2011; Handcock, 2003). Moreover, the estimation process, when based on Markov Chain Monte Carlo (MCMC) methods, often fails due to the complexity and size of the data. Larger networks may imply degenerate models with poor sampling properties and longer mixing times (Bhamidi et al., 2011).
@@
Throughout this paper, we consider a signed network with $N\in\mathbb{N}$ nodes. The adjacency matrix for this network is denoted as $\mathbf{y}=(y_{i,j})\in\mathcal{S}^{N\times N}\coloneqq\mathcal{Y}$, with $\mathcal{S}\coloneqq\{\!``-",``0",``+"\!\}$, where $\mathcal{Y}$ is the set of all observable signed adjacency matrices among $N$ fixed nodes. Therefore, $y_{i,j}=``-"$ indicates a negative edge, $y_{i,j}=``+"$ represents a positive edge, and $y_{i,j}=``0"$ signifies the absence of an edge between nodes $i$ and $j$. In the undirected case, we have $y_{i,j}=y_{j,i}$, implying that the adjacency matrix is symmetric. For this paper, we assume that the network is undirected and has no self-loops, so $y_{i,i}=``0"$. However, extensions to the directed case follow naturally. By $\mathbf{y}_{(-ij)}$, we denote $\mathbf{y}$ excluding $y_{i,j}$. The set of nodes is partitioned into $K\in\mathbb{N}$ disjoint blocks. Let $N_{k}\in\mathbb{N}$ denote the number of nodes in block $k$, for $k=1,\ldots,K$, such that $\sum_{k=1}^{K}N_{k}=N$. The unobservable vector $\bm{z}_{i}=(z_{i1},\dots,z_{iK})$ denotes the block membership of each node, where $z_{i,k}=1$ if node $i$ belongs to block $k$, and $z_{i,k}=0$ otherwise. Let $\mathbf{y}_{k,l}\in\mathcal{S}^{N_{k}\times N_{l}}$ denote the submatrix of $\mathbf{y}$ that contains all edges between nodes in block $k$ and nodes in block $l$, for $k,l=1,\ldots,K$. Throughout this paper, capital letters refer to random variables, lowercase letters to their realizations, and bold letters denote vectors or matrices.
@@ 2 Existing Models for Signed Networks
The Signed Exponential Random Graph Model (SERGM, Fritz et al., 2025) posits that an observed network can be fully characterized through a set of sufficient statistics, which capture both endogenous dependencies, which are structural features of the network, and exogenous influences, relating to information such as covariate effects independent of the network (Lusher et al., 2013). Fritz et al. (2025) defined the probability of observing $\mathbf{y}\in\mathcal{Y}$ is given by
@@
where $\pi_{k,l}(y)$ with $y\in\mathcal{S}$ denotes the probability of $Y_{i,j}=y$ with $\sum_{y\,\in\,\mathcal{Y}_{i,j}}\pi_{k,l}(y)=1$ and ${\bm{\pi}}=(\pi_{k,l}(y))$ (Li et al., 2023). A shortcoming of this model is its simplicity: The probability to observe a negative, neutral, or positive edge between nodes $i$ and $j$ solely depends on the block memberships of the involved nodes. This dependence structure is not able to capture foundational theories about signed networks, such as the structural balance theory.
@@ 3 Exponential Random Graph Model for Signed Networks under Local Dependence
where $\bm{\theta}\coloneqq\operatorname{vec}(\bm{\theta}_{1,1},\ldots,\bm{\theta}_{K,K},\bm{\theta}_{1,2},\ldots,\bm{\theta}_{K-1,K})$ denotes the vector of both between-block and within-block parameters with $\operatorname{vec}$ defining a function that stacks its arguments vertically. Unlike with stochastic block models, this model does not assume that edges within and between blocks are independent. Rather, only the independence of between block edges is guaranteed, while within block edges can be strongly dependent.
@@
where $\bm{\theta}_{k,k}\in\mathbb{R}^{p}$ denotes the vector of within-block parameters for block $k$, and $\bm{s}:\mathcal{Y}_{k,k}\rightarrow\mathbb{R}^{p}$ is a function calculating the sufficient statistics. This function includes both dyad-independent and dyad-dependent variables, since we assume that nodes are aware of the activities within their block. The number of sufficient statistics used to model within-block networks $p_{w}\in\mathbb{N}$ is assumed to be constant across blocks.
@@
The model would constitute a curved ERGM with local dependence if the relationship between the population-level coefficients and the block-specific covariates in equation (5) were non-linear (Hunter and and Handcock, 2006). Block-specific covariates $\bm{v}_{k}$, such as block size $N_{k}$ or $\log(N_{k})$, further enable size-dependent parametrizations (see, e.g., Butts and and Almquist, 2015 and Krivitsky et al., 2011). Larger blocks will typically have a lower density, however using non-size dependent parameters would lead to preservation of the density. Let $\beta_{w,a,b}$, $\theta_{k,k,a}$ , and $v_{k,b}$ denote the $(a,b)$th entry of $\bm{\beta}_{w}$, the within-block effect of the $a$th statistic in $\bm{\theta}_{k,k}$, and value of the $b$th entry of $\bm{v}_{i}$, respectively. Then $\theta_{k,k,b}$ is given by $\beta_{w,a,b}\,\bm{v}_{k}$ and $\beta_{w,a,b}$ can be interpreted as the increase of $\theta_{k,k,b}$ when $v_{k,b}$ is increase by one unit. Covariates $\bm{u}_{k,l}$ typically only include a constant, dummy variables for specific pairs of blocks, or functions of $\bm{v}_{k}$ and $\bm{v}_{l}$.
@@
Thus, $\theta_{k,l,+}$ and $\theta_{k,l,-}$ represent log-odds of observing a positive and negative edge between blocks $k$ and $l$ compared to observing no edge. In this form, each $\bm{s}(\mathbf{y}_{k,k})$ and $\bm{h}(\mathbf{y}_{k,l})$ acts like an intercept term: it aggregates the number of edges of a particular sign, and its corresponding parameter controls the baseline probability of that edge type appearing between nodes with the given block memberships. For the within-block model, we can also represent this in the linear predictor structure from Equation (5). In this case each $\bm{v}_{k}\in\mathbb{R}^{K}$ is a one-hot vector with a 1 in the $k$th position and 0 elsewhere. Similarly, for the between-block model, $\bm{u}_{k,l}\in\mathbb{R}^{K}$ is a vector with 1s in position $k$ and $l$. To reduce model complexity and avoid estimating block-pair-specific parameters, we assume homogeneous parameters across all between-block pairs. In this case, we define $\bm{u}_{k,l}=1$, so al $\bm{\theta}_{k,l}=\bm{\beta}_{b}\in\mathbb{R}^{1\times 2}$ In the following examples, this parameter vector remains fixed, as no dyad-independent terms are added to the between-block model.
@@
To capture more complex local structures beyond count variables such as edge and triad counts, we include geometrically weighted positive/negative degree and edgewise shared partner statistics, specifically the geometrically weighted “enemy of my enemy is my friend” triad. These geometrically weighted statistics reduce the risk of model degeneracy by reducing the weight of high degree nodes or high shared partner counts, thereby improving model stability. The decay parameter $\omega$ controls this discounting: smaller values of $\omega$ place more weight on higher-degree nodes or triads with many shared partners, while larger values of $\omega$ reduce their influence. For a more detailed overview of geometrically weighted statistics, see Hunter (2007).
Define the geometrically weighted positive degree statistic as:
@@
counts nodes in block $k$ with positive or negative degree exactly $d$.
Define the geometrically weighted edgewise shared partner statistic as:
@@
Here, $\text{ESE}_{d}^{y}$ counts edges of sign $y$ sharing exactly $d$ common enemies, and $\text{ESF}_{d}^{y}$ counts edges of sign $y$ sharing exactly $d$ common friends.
The vector of sufficient statistics for the within-block model is then defined as:
@@ 4 Scalable Estimation
following the practice for SSBMs (Li et al., 2023). Combining (6) with the conditional model $\mathbf{Y}\mid\bm{Z}=\bm{z}$ (defined in (3)) yields a latent variable model (see, David et al., 2011 for an introduction to the field). The log-likelihood for parameters $\bm{\theta}$ and $\gamma$ is
@@
The optimal solution in Step 1 is $A(\bm{Z})=\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$, which does not admit a closed-form expression (Matias and Robin, 2014). For SBMs of moderate size, Gibbs sampling based on the full conditional distributions $Z_{i}\mid Z_{1},...,Z_{i-1},Z_{i+1},...,Z_{N}$ can be employed to approximate $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$ (Nowicki and and Snijders, 2001). However, for larger networks, as considered in this paper, variational methods offer scalable approximations. For this approach, we restrict $A$ to a tractable family of distributions parametrized by $\bm{\alpha}$ and approximate $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})\approx A_{\bm{\alpha}}(\bm{z})$, where $A_{\bm{\alpha}}(\bm{z})$ minimizes the Kullback-Leibler divergence to $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$. An explicit form of this variational family is provided in Section 4.2.
Still, the dependence of within-block networks (4) inhibits adaption to our setting. In Section 4.1, we adapt the result of Babkin et al. (2020) to signed networks showing that under some conditions the probability (3) can be approximated by the probability of a SSBM, mentioned in Section 2 and Example 1. This result justifies the use scalable optimization methods originally developed for SBMs. More concretely, in Section 4.2 we adapt the approach of Vu et al. (2013) based on MM Steps to the signed setting. After convergence, a point estimate $\hat{\bm{z}}$ is derived Section 4.3 discusses the estimation of $\bm{\theta}$ conditional on $\bm{Z}=\hat{\bm{z}}$. Section 4.4 proposes methods to incorporate uncertainty in $\mathbf{Z}$ neglected by this deterministic assignment.

### v2

@@ Abstract
Traditional network analysis focuses on binary edges, while real-world relationships are more nuanced, encompassing cooperation, neutrality, and conflict. The rise of negative edges in social media discussions spurred interest in analyzing signed interactions, especially in polarized debates. However, the vast data generated by digital networks presents challenges for traditional methods like Stochastic Block Models (SBM) and Exponential Family Random Graph Models (ERGM), particularly due to the homogeneity assumption and global dependence, which become increasingly unrealistic as network size grows. To address this, we propose a novel method that combines the strengths of SBM and ERGM while mitigating their weaknesses by incorporating local dependence based on nonoverlapping blocks. Our approach involves a two-step process: first, decomposing the network into sub-networks using SBM approximation, and then estimating parameters using ERGM methods. We validate our method on large synthetic networks and apply it to a signed Wikipedia network of thousands of editors. Through the use of local dependence, we find patterns consistent with structural balance theory.
@@ 1 Introduction
Nevertheless, the modeling of signed networks remains challenging. On the one hand, the class of models specifically designed for signed networks is relatively limited, as most models were developed for unsigned networks. On the other hand, the massive size of modern networks poses separate challenges, as even for unsigned networks many models struggle to scale computationally. At the intersection of these two issues, the need to model both signed relations and large networks, there are little to no methods available.
Two of the most prominent models for network analysis are Stochastic Block Models (SBM, Nowicki and Snijders, 2001) and Exponential Family Random Graph Models (ERGM, Lusher et al., 2013). An SBM is a latent variable model in which all dependence between edges is accounted for by latent group membership. ERGMs, on the other hand, allow for more complex local dependencies through so-called sufficient statistics (De Nicola et al., 2023). Signed network extensions exist for both SBM and ERGM (Jiang, 2015; Fritz et al., 2025) with inherent shortcomings: While the SBM can represent cohesive subgroups among the nodes, it implies conditional independence for forming edges between nodes. This assumption is unrealistic for networks exhibiting more dependence among edges. ERGMs, on the other hand, offer a flexible framework to analyze general network properties. For a parsimonious model, these properties are commonly assumed to be homogeneous across the entire network. The induced global dependence structure, where each edge may technically depend on every other edge in the same way, gives rise to undesirable behavior specifically for large networks (Schweinberger, 2011; Handcock, 2003). Moreover, the estimation process, when based on Markov Chain Monte Carlo (MCMC) methods, often fails due to the complexity and size of the data. Larger networks may imply degenerate models with poor sampling properties and longer mixing times (Bhamidi et al., 2011).
@@
Throughout this paper, we consider a signed network with $N\in\{2,\ldots\}$ nodes. The adjacency matrix for this network is denoted as $\mathbf{y}=(y_{i,j})\in\mathcal{S}^{N\times N}\coloneqq\mathcal{Y}$, with $\mathcal{S}\coloneqq\{\!``-",``0",``+"\!\}$. The set of all observable signed adjacency matrices among $N$ fixed nodes is defined by $\mathcal{Y}$. Therefore, $y_{i,j}=``-"$ indicates a negative edge, $y_{i,j}=``+"$ represents a positive edge, and $y_{i,j}=``0"$ signifies the absence of an edge between nodes $i$ and $j$. In the undirected case, we have $y_{i,j}=y_{j,i}$, implying that the adjacency matrix is symmetric. For this paper, we assume that the network is undirected and has no self-loops, so $y_{i,i}=``0"$. However, extensions to the directed case follow naturally. By $\mathbf{y}_{(-ij)}$, we denote $\mathbf{y}$ excluding $y_{i,j}$. The set of nodes is partitioned into $K\in\{2,\ldots\}$ disjoint blocks. Let $N_{k}\in\{2,\ldots\}$ denote the number of nodes in block $k$, for $k=1,\ldots,K$, such that $\sum_{k=1}^{K}N_{k}=N$. The unobservable vector $\bm{z}_{i}=(z_{i1},\dots,z_{iK})$ denotes the block membership of each node, where $z_{i,k}=1$ if node $i$ belongs to block $k$, and $z_{i,k}=0$ otherwise. Let $\mathbf{y}_{k,l}\in\mathcal{S}^{N_{k}\times N_{l}}$ denote the submatrix of $\mathbf{y}$ that contains all edges between nodes in block $k$ and nodes in block $l$, for $k,l=1,\ldots,K$. Throughout this paper, capital letters refer to random variables, lowercase letters to their realizations, and bold letters denote vectors or matrices.
@@ 2 Existing Models for Signed Networks
The Signed Exponential Random Graph Model (SERGM, Fritz et al., 2025) posits that an observed network can be fully characterized through a set of sufficient statistics, which capture both endogenous dependencies, which are structural features of the network, and exogenous influences, relating to information such as covariate effects independent of the network (Lusher et al., 2013). Fritz et al. (2025) defined the probability of observing $\mathbf{y}\in\mathcal{Y}$ by
@@
where $\pi_{k,l}(y)$ with $y\in\mathcal{S}$ denotes the probability of $Y_{i,j}=y$ with $\sum_{y\,\in\,\mathcal{Y}_{i,j}}\pi_{k,l}(y)=1$ and ${\bm{\pi}}=(\pi_{k,l}(y))$ (Li et al., 2023). A shortcoming of this model is its simplicity: The probability to observe a negative, neutral, or positive edge between nodes $i$ and $j$ solely depends on the block memberships of the involved nodes. This dependence structure is not able to capture foundational theories about signed networks, such as the structural balance theory (Heider, 1946).
@@ 3 Exponential Random Graph Model for Signed Networks under Local Dependence
where $\bm{\theta}\coloneqq\operatorname{vec}(\bm{\theta}_{1,1},\ldots,\bm{\theta}_{K,K},\bm{\theta}_{1,2},\ldots,\bm{\theta}_{K-1,K})$ denotes the vector of both between-block and within-block parameters with $\operatorname{vec}$ defining a function that stacks its arguments vertically. Unlike with stochastic block models, this model does not assume that edges within and between blocks are independent. Rather, only the independence of between-block edges is guaranteed, while within-block edges can be strongly dependent.
@@
where $\bm{\theta}_{k,k}\in\mathbb{R}^{p}$ denotes the vector of within-block parameters for block $k$, and $\bm{s}:\mathcal{Y}_{k,k}\rightarrow\mathbb{R}^{p}$ is a function calculating the sufficient statistics. This function includes both dyad-independent and dyad-dependent variables, since we assume that nodes are aware of the activities within their block. The number of sufficient statistics used to model within-block networks $p\in\{1,\ldots\}$ is assumed to be constant across blocks.
@@
The model would constitute a curved ERGM with local dependence if the relationship between the population-level coefficients and the block-specific covariates in equation (5) were non-linear (Hunter and Handcock, 2006). Block-specific covariates $\bm{v}_{k}$, such as block size $N_{k}$ or $\log(N_{k})$, further enable size-dependent parametrizations (see, e.g., Butts and Almquist, 2015 and Krivitsky et al., 2011). Larger blocks will typically have a lower density, however using non-size dependent parameters would lead to preservation of the density. Let $\beta_{w,a,b}$, $\theta_{k,k,a}$ , and $v_{k,b}$ denote the $(a,b)$th entry of $\bm{\beta}_{w}$, the within-block effect of the $a$th statistic in $\bm{\theta}_{k,k}$, and value of the $b$th entry of $\bm{v}_{i}$, respectively. Then $\theta_{k,k,b}$ is given by $\beta_{w,a,b}\,\bm{v}_{k}$ and $\beta_{w,a,b}$ can be interpreted as the increase of $\theta_{k,k,b}$ when $v_{k,b}$ is increased by one unit. Covariates $\bm{u}_{k,l}$ typically only include a constant, dummy variables for specific pairs of blocks, or functions of $\bm{v}_{k}$ and $\bm{v}_{l}$.
@@
Thus, $\theta_{k,l,+}$ and $\theta_{k,l,-}$ represent log-odds of observing a positive and negative edge between blocks $k$ and $l$ compared to observing no edge. In this form, each $\bm{s}(\mathbf{y}_{k,k})$ and $\bm{h}(\mathbf{y}_{k,l})$ acts like an intercept term: it aggregates the number of edges of a particular sign, and its corresponding parameter controls the baseline probability of that edge type appearing between nodes with the given block memberships. For the within-block model, we can also represent this in the linear predictor structure from Equation (5). In this case, each $\bm{v}_{k}\in\mathbb{R}^{K}$ is a one-hot vector with a 1 in the $k$th position and 0 elsewhere. Similarly, for the between-block model, $\bm{u}_{k,l}\in\mathbb{R}^{K}$ is a vector with 1s in position $k$ and $l$. To reduce model complexity and avoid estimating block-pair-specific parameters, we assume homogeneous parameters across all between-block pairs. In this case, we define $\bm{u}_{k,l}=1$, so all $\bm{\theta}_{k,l}=\bm{\beta}_{b}\in\mathbb{R}^{1\times 2}$ In the following examples, this parameter vector remains fixed, as no dyad-independent terms are added to the between-block model.
@@
To capture complex local structures beyond simple count statistics such as edge or triad counts, we incorporate geometrically weighted degree and edgewise shared partner statistics, including a geometrically weighted version of the “enemy of my enemy is my friend” triad configuration. These statistics account for variation in local connectivity patterns, particularly in networks characterized by degree heterogeneity and a tendency toward triadic closure. By down-weighting high-degree nodes and configurations with many shared partners or degrees, geometric weighting reduces the influence of extreme values and improves the stability of parameter estimation. The decay parameter $\omega$ controls this down-weighting: smaller values of $\omega$ place more weight on high degree nodes or triads with many shared partners, while larger values of $\omega$ reduce their influence. A more detailed overview of geometrically weighted statistics is provided in Hunter (2007).
The geometrically weighted degree statistic for positive or negative edges within block $k$ is defined as:
@@
counts nodes in block $k$ with a positive or negative degree of $d$. Analogously, the geometrically weighted edgewise shared partner statistic is given by:
@@
Here, $\text{ESE}_{d}^{y}$ counts edges of sign $y$ sharing exactly $d$ common enemies, and $\text{ESF}_{d}^{y}$ counts edges of sign $y$ sharing exactly $d$ common friends. The vector of sufficient statistics for the within-block model is then defined as:
@@ 4 Scalable Estimation
following the practice for SSBMs (Li et al., 2023). Combining (6) with the conditional model $\mathbf{Y}\mid\bm{Z}=\bm{z}$ (defined in (3)) yields a latent variable model (see, David et al., 2011, for an introduction to the field). The log-likelihood for parameters $\bm{\theta}$ and $\gamma$ is
@@
The optimal solution in Step 1 is $A(\bm{Z})=\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$, which does not admit a closed-form expression (Matias and Robin, 2014). For SBMs of moderate size, Gibbs sampling based on the full conditional distributions $Z_{i}\mid Z_{1},...,Z_{i-1},Z_{i+1},...,Z_{N}$ can be employed to approximate $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$ (Nowicki and Snijders, 2001). However, for larger networks, as considered in this paper, variational methods offer scalable approximations. For this approach, we restrict $A$ to a tractable family of distributions parametrized by $\bm{\alpha}$ and approximate $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})\approx A_{\bm{\alpha}}(\bm{z})$, where $A_{\bm{\alpha}}(\bm{z})$ minimizes the Kullback-Leibler divergence to $\mathbb{P}(\bm{Z}=\bm{z}\mid\mathbf{Y}=\mathbf{y})$. An explicit form of this variational family is provided in Section 4.2.
Still, the dependence of within-block networks (4) inhibits adaption to our setting. In Section 4.1, we adapt the result of Babkin et al. (2020) to signed networks showing that under some conditions the probability (3) can be approximated by the probability of an SSBM, mentioned in Section 2 and Example 1. This result justifies the use of scalable optimization methods originally developed for SBMs. More concretely, in Section 4.2 we adapt the approach of Vu et al. (2013) based on MM Steps to the signed setting. After convergence, a point estimate $\hat{\bm{z}}$ is derived. Section 4.3 discusses the estimation of $\bm{\theta}$ conditional on $\bm{Z}=\hat{\bm{z}}$. Section 4.4 proposes methods to incorporate uncertainty in $\mathbf{Z}$ neglected by this deterministic assignment.